# myboxenhanced,colback=blue!5!white,colframe=blue!75!black,boxrule=0.6mm,fontupper=

### v1

@@ 3 Side Information and its Usage
In addition to the characterization of the problem by an MDP, there can still be additional information that could potentially improve performance on additional metrics such as Sample Efficiency, Generalization, Interpretability, and Safety. We call this Side Information (also referred to as privileged information). For the (semi-) supervised and unsupervised settings, side information is any additional information $z\in\mathcal{Z}$ that, while being neither part of the input nor the output space, can potentially contribute to the learning process (?).
Translated to the RL setting, this can be understood as additional information $z$ not provided in the original MDP definition $\mathcal{M}$. Such information can be incorporated into the learning process by a function $\beta:\mathcal{Z}\to\mathcal{X}$, where $\mathcal{X}\in\Omega$. Thus, side information can be incorporated into the RL pipeline by biasing one or more of the components shown in Figure 2.
The natural follow-up question, then, becomes the impact of incorporating side information into the learning pipeline. In this work, we specifically focus on four ways in which side information can be used and formally define them in the following sections.
@@ 3.1 Sample Efficiency
Sample Efficiency is intimately tied to the idea of the Sample Complexity of RL. To formally define it, we use the notion of the Sample Complexity of Exploration (?): Given fixed parameters $\epsilon,\delta>0$, if the difference between the value functions of a learned policy $\pi$ and an optimal policy $\pi^{*}$ is $||Q^{\pi}(s,a)-Q^{\star}(s,a)||_{\infty}>\epsilon$, then we call this a mistake. If, with a probability of $1-\delta$, the total number of mistakes at a timestep $t$ is $\zeta(\epsilon,\delta)$, then we call $\zeta$ the sample complexity of exploration.
Incorporating side information leads to a reduction in $\zeta$, thus, improving the sample efficiency. Intuitively, if a pipeline demonstrates a higher reward than a baseline for the same number of timesteps, then we consider it more sample efficient. However, methods can additionally make specific claims on $\zeta$ by utilizing certain assumptions about the problem itself (?, ?, ?).
@@
One specific way to improve the sample complexity of exploration is to impact the exploration mechanism using side information directly. ? (?) categorize exploration methods based on the type of information that an agent uses to explore the world into the following categories: (i) Reward-Free Explorationmethods in which extrinsic rewards do not affect the choice of action. Instead, they rely on intrinsic forms of exploration. (ii) Randomized Action Selectionmethods use estimated value functions, policies, or rewards to induce exploratory behavior. (iii) Optimism/Bonus-Based Explorationmethods use the optimism in the face of uncertainty paradigm to prefer actions with higher uncertain values. (iv) Deliberate Explorationmethods that either use posterior distributions over dynamics (Bayesian setup) or meta-learning techniques to optimally solve exploration and (v) Probability Matchingmethods that use heuristics to select the next action. Incorporating side information into any of these methods generally leads to better sample efficiency, particularly by improving the state-space coverage of the exploration mechanism.
@@ 3.2 Transfer and Generalization
Transfer and generalization encompass performance metrics that measure how an RL agent performs on a set of different MDPs: Transfer evaluates how well an agent, trained on some MDP $\mathcal{M}_{i}$ performs on some other MDP $\mathcal{M}_{j}$. This can be either done in a zero-shot manner, where the agent is not fine-tuned on $\mathcal{M}_{j}$, or in a few-shot manner, where the agent gets to make some policy updates on $\mathcal{M}_{j}$ to learn as fast as possible. Generally, the performance gap between the two MDPs determines the transfer performance.
@@ 3.3 Interpretability
Interpretability refers to a mechanistic understanding of a system to make it more transparent. ? (?) enumerate three fundamental properties of model interpretability: (i) Simulatabilityrefers to the ability of a human to simulate the inner working of a system, (ii) Decomposabiltiyrefers to adding intuitive understanding to individual working parts of a system, (iii) Transparencyrefers to improving the understanding of a system’s function (e.g., quantifying its convergence properties).
Given the coupled nature of individual parts of an RL pipeline, adding interpretability amounts to being able to learn a policy for the MDP that adheres to one of multiple such properties. Incorporating side information can help take steps towards improved performance on all three of these metrics, depending on the nature of side information and what it encompasses. We consider claims on interpretability based on whether a given work additionally addresses at least two of these metrics.
@@ 3.4 Safety
Safety is the idea of learning policies that maximize the expectation of the return in problems in which it is important to ensure reasonable system performance and respect safety constraints during the learning and or deployment processes.
While Safety in RL is a vast field in and of itself (?), we consider two specific categories in this work: Safe Learning with constraints and Safe Exploration. The former modifies the expected return account for one or more constraints $c_{i}\in C$, and the general form can be written as
$\max_{\pi\in\Pi}\mathbb{E}_{\pi}(G)\,\,\,s.t.\,\,\,c_{i}=\{h_{i}\leq\alpha\}$ (6)
where $h_{i}$ is a function related to the return and $\alpha$ is the threshold restricting the values of this function. Consequently, side information can be used in the formulation of such constraints.
On the other hand, Safe Exploration modifies the exploration process subject to external knowledge, which in our case translates to incorporating side information into the exploration process. While intuitively, this overlaps with the usage of side information for directed exploration, a distinguishing feature of this work is the final goal of this directed exploration to be safety, which might potentially come at the cost of sample efficiency and/or generalization.
@@ 4 Structure as Side Information
Structure can be considered as a particular kind of side information available to a learning agent. To build an intuition about what we mean by this, consider the task of managing a large factory with many production cells (example taken from ? (?)). If a cell positioned early in the production line generates faulty parts, the whole factory may be affected. However, the quality of the parts a cell generates depends directly only on the state of this cell and the quality of the parts it receives from neighboring cells. Additionally, the cost of running the factory depends, among other things, on the sum of the costs of maintaining each local cell. Finally, while a cell responsible for anodization may receive parts directly from any other cell in the factory, a work order for a cylindrical part may restrict this dependency to cells with a lathe. In the context of producing cylindrical parts, the quality of the anodized parts depends directly only on the state of cells with a lathe. Thus, by incorporating information about the additive nature of production, costs, and the context of the part that needs to be produced, the learning pipeline can be imbued with better objectives such as improved sample efficiency or robustness of a learned policy to changing factory conditions.
The vanilla MDP framework does not require incorporating such additive information into the learning pipeline. Thus, biasing the pipeline with this amounts to incorporating side information. However, this information particularly helps decompose the complicated learning problem into additive sub-parts that can be learned independently and, potentially, more efficiently. Thus, structure is a particular kind of side information that provides the learning pipeline with knowledge about decomposability in the learning problem.
In this section, we discuss the relationship between structure and decomposability. In Section 4.1, we explain the impact of structural side information by explaining how it decomposes complex systems and categorizes such decompositions into four archetypes. In Section 4.2 - Section 4.5, we then discuss each of these archetypes further to connect them with existing literature.
@@ 4.1 Decomposability and Structural Archetypes
Decomposability is the property of a system that allows breaking it down into smaller components or subsystems that can be analyzed, understood, and potentially learned more efficiently than the larger system, independently (?). In a decomposable system, the short-term behavior of each subsystem is approximately independent of the short-term behavior of the other subsystems. In the long run, the behavior of any one subsystem depends on the behavior of the other subsystems only in an aggregated way.
Concerning the RL pipeline in Figure 2, we can see decomposability along two axes: (i) Problem Decompositioni.e., the environment parameterization, states, actions, transitions, and rewards; (ii) Solution Decompositioni.e., the learned policies, value functions, models, and training procedures. The spectrum of decomposability (?) provides an intuitive way to understand where a system lies in this regard. On one end of the spectrum, there are problems that are non-decomposable, while on the other end, problems can be decomposed into weakly interacting sub-problems. Consequently, solutions on the former end are monolithic, while those on the latter end are modular. We capture this problem-solution interplay by marking four different archetypes of decomposability, as shown in Figure 3.
Figure 3: Spectrum of Decomposability and Structural Archetypes. On the left end of the spectrum exist monolithic structural decompositions where knowledge about a latent subspace of $\mathcal{X}$ can be learned and incorporated as an inductive bias. As we move towards the right, we see the possibility of learning multiple independent subspaces, albeit in a monolithic solution. These are factored decompositions. Further ahead, we see the emergence of interactionally complex decompositions, where knowledge about factorization and how they relate to each other can be incorporated into the learning process. We call these relational decompositions. Finally, we see fully distributed subsystems that can be incorporated and learned using individual policies. We call these modular decompositions.
@@ 4.2 Latent Decomposition
Latent decompositions are monolithic and can be useful in complex environments where the underlying structure is either unclear or non-stationary. Under this view, a pipeline component $\mathcal{X}$ can be approximated by a latent representation $\kappa$, which can then be integrated into the learning process. The quantities in the update that depend on $\mathcal{X}$ can now be re-conditioned on $\kappa$, which helps in improving performance by reducing the dimensionality of $\mathcal{X}$.
@@
Latent state spaces have been classically explored through the Latent MDP literature (?), where the aim is to discover a latent representation of the state space that is sufficient to learn an optimal policy. Latent states are used for tackling rich observation spaces, where Block MDPs (?) and Contextual MDPs (?) have shown success in generalization problems. Latent actions have been similarly explored in settings with stochastic action sets (?).
Given an encoder $\phi:\mathcal{S}\times\mathcal{A}\to\mathcal{\kappa}$ and a decoder $\beta:\mathcal{\kappa}\to\mathcal{S}$, the latent state-action formulation allows decomposing transition dynamics into a low-rank approximation,
$P(s^{\prime}\mid s,a)=\langle\phi(\kappa\mid s,a)\beta(s^{\prime}\mid\kappa)\rangle.$
@@
While latent states allow decomposing transition matrices, another way to approach the problem directly is to decompose transition matrices into low-rank approximations:
$P(s^{\prime}\mid s,a)=\phi(s^{\prime}\mid s,a)\beta(s^{\prime}\mid s,a).$
Linear MDPs (?) and corresponding applications in Model-based RL (?, ?) have studied this form of direct decomposition.
A similar decomposition can be applied to rewards as well. ? (?) have primarily explored this in noisy reward settings where the reward signal is assumed to be generated from a latent function that can be learned as an auxiliary learning objective.
@@ 4.3 Factored Decomposition
The factored decomposition moves slightly away from the monolithic nature by decomposing $\mathcal{X}$ into (latent) factors $\kappa_{1},\dots,\kappa_{n}$. Thus, the spaces become inner products of the individual factor spaces. A crucial aspect that differentiates factorization is that the factors can potentially impose conditional independence in their effects on the learning dynamics.
@@
Factored state and action spaces have been explored in the Factored MDPs (?, ?, ?). Methods in this setting traditionally capture next state distribution using mechanisms such as Dynamic Bayesian Networks (?).
Factorization in the action space has also been used for tackling high-dimensional actions (?). These methods either impose a factorized structure on subsets of a high-dimensional action set (?) or impose this structure through the Q-values that lead to the final action (?). Crucially, these methods can potentially exploit some form of independence resulting from such factorization, either in the state representations or transitions.
@@
Combined with factored states or modeled independently, factored rewards have been used to model perturbed rewards and multi-objective settings (?). While Factored MDPs do not naturally lead to factored policies, combining state and reward factorization can lead to factorization of Q-values (?, ?).
@@ 4.4 Relational Decomposition
Relational decompositions add a further notion of separability where in addition to the factored subsets, capturing immutable relations between them becomes important as well. Usually, these relations exist between entities in a scene and are used to formulate learning methods based on inductive logic (?). Traditionally, such relations were limited to first-order logic, but the relational structure has also been captured through graphs.
The relational assumption posits that a space of predicates $\beta$ can ground these entities, and it can be modeled as a set of rules (such as inductive logic) that define how $z_{i},z_{j}\in Z$ interact with each other. An extension of this is by capturing $\beta$ as a graph $G=\langle V,E\rangle$ where the vertices are $z_{i}\in Z$ and the edges represent the relationship between them.
Using such a representation allows us to talk about generalization over the entities $\kappa_{i},\kappa_{j}$, and forms of $\beta$. This helps us circumvent the dimensionality of enumerative spaces.
Relational States and Actions
Classically, relational representations have been used to model state spaces in Relational MDPs (?) and Object-Oriented MDPs (?, ?). They represent factored state spaces using first-order representations consisting of objects, predicates, and functions to describe a set of ground MDPs. Such representations can capture interactionally complex structure that exists between entities much more efficiently. Additionally, permutations of the interactions between the entities can help define new MDPs that differ in their dynamics, thus, contributing towards work in generalization.
@@
Action relations help tackle instances where the agent has multiple possible actions available and the set of actions is significantly large. These methods capture relations using either attention mechanisms (?, ?) or graphs (?), thus offering scalability to high-dimensional action spaces.
Relational Value Functions and Policies
Traditional work in Relational MDPs has also explored ways to represent and construct first-order representations of value functions and/or policies to generalize to new instances. These include Regression Trees (?), Decision Lists (?), Algebraic Decision Diagrams (?), and Linear Basis Functions (?, ?). Recent approaches have started looking into DNN representations (?, ?), with extensions into modeling problem aspects such as morphology in Robotic tasks (?) in a relational manner.
Relational Tasks

### v3

@@ 4 Side Information and its Usage
In addition to the characterization of the problem by an MDP, there can still be additional information that could potentially improve performance on additional metrics such as Sample Efficiency, Generalization, Interpretability, and Safety. We call this Side Information (also called privileged information). For the (semi-) supervised and unsupervised settings, side information is any additional information that, while neither part of the input nor the output space, can potentially contribute to the learning process (?).
Translated to the RL setting, this can be understood as additional information not provided in the original MDP definition $\mathcal{M}$. Side information can be incorporated into the RL pipeline by biasing one or more components shown in Figure 2. Mathematically, we can express this using some function $\beta$ that conditions the pipeline on side information by augmenting it with a function $Z$.
$\beta:\Omega\to\Omega\times\mathcal{Z}$
This implies that we now augment our tuple $\Omega$ with an additional function $Z$ that operates on other tuple elements $\mathcal{X}\in\Omega$. For example, incorporating side information could be used to learn state abstractions by adding an encoder $Z$ to map the state space $\mathcal{S}$ to a latent representation $\kappa$ that can be used for control. We discuss the general templates for $Z$ in Section 5 and classify different methods of biasing $\Omega$ with $Z$ into patterns in Section 6. The natural follow-up question, then, becomes the impact of incorporating side information into the learning pipeline. In this work, we focus on four ways side information can be used and formally define them in the following sections.
@@ 4.1 Sample Efficiency
Sample Efficiency is intimately tied to the Sample Complexity of RL. Intuitively, if a pipeline demonstrates a higher reward than a baseline for the same number of timesteps, we consider it more sample-efficient. To formally define it, we use the notion of the Sample Complexity of Exploration (?, ?): Given some $\epsilon>0$, we can define the sample complexity as the number of timesteps $t$ after which the policy produces a value $V^{\pi}<V^{\star}-\epsilon$. This definition by ? (?) directly measures the number of times the agent acts poorly (quantified by $\epsilon$) and views “fast” learners as those that act poorly as few times as possible. Incorporating side information leads to a reduction in sample complexity, thus improving the sample efficiency.
@@
One specific way to improve the sample complexity of exploration is to directly impact the exploration mechanism using side information. ? (?) categorize exploration methods based on the type of information that an agent uses to explore the world into the following categories: (i) Reward-Free Explorationmethods in which extrinsic rewards do not affect the choice of action. Instead, they rely on intrinsically motivated forms of exploration, such as diversity maximization (?). (ii) Randomized Action Selectionmethods use estimated value functions, policies, or rewards to induce exploratory behavior. (iii) Optimism/Bonus-Based Explorationmethods use the optimism in the face of uncertainty paradigm to prefer actions with higher uncertain values. (iv) Deliberate Explorationmethods that either use posterior distributions over dynamics (Bayesian setup) or meta-learning techniques to optimally solve exploration and (v) Probability Matchingmethods that use heuristics to select the next action. Incorporating side information into any of these methods generally improved the state-space coverage of the exploration mechanism. We specifically cover methods that impact the exploration mechanism to improve sample efficiency and/or generalization.
@@ 4.2 Transfer and Generalization
Transfer and generalization encompass performance metrics that measure how an RL agent performs on a set of different MDPs: Transfer evaluates how well an agent, trained on some MDP $\mathcal{M}_{i}$, performs on another MDP $\mathcal{M}_{j}$. This can be either done in a zero-shot manner, where the agent is not fine-tuned on $\mathcal{M}_{j}$, or in a few-shot manner, where the agent gets to make some policy updates on $\mathcal{M}_{j}$ to learn as fast as possible. Generally, the performance gap between the two MDPs determines the transfer performance.
@@
Our patterns framework in Section 6 proposes a new way to categorize these approaches into design patterns. Consequently, we cover all such forms of handling generalization.
@@ 4.3 Interpretability
Interpretability refers to a mechanistic understanding of a system to make it more transparent. ? (?) enumerate three fundamental properties of model interpretability: (i) Simulatabilityrefers to the ability of a human to simulate the inner workings of a system, (ii) Decomposabilityrefers to adding intuitive understanding to individual working parts of a system, (iii) Transparencyrefers to improving the understanding of a system’s function (such as quantifying its convergence properties).
Given the coupled nature of individual parts of an RL pipeline, adding interpretability amounts to learning a policy for the MDP that adheres to at least one of multiple such properties. Incorporating side information can help improve performance in all three aspects, depending on the nature of side information and what it encompasses. However, we do not explicitly provide a formal metric for interpretability due to the potentially subjective nature of such metrics, especially in the case of RL, where performances on such metrics might differ depending on the environment. We instead look for interpretability through the lens of decomposability, which we discuss further in Section 5, by specifically checking for whether the decompositions leveraged by the methods are individually simulatable or add transparency to the action-selection mechanism.
@@ 4.4 Safety
Safety refers to learning policies that maximize the expectation of the return in problems in which it is important to ensure reasonable system performance and/or respect safety-related constraints during the learning and/or deployment processes. For example, Model-based RL methods usually learn a model of the environment and then use it to plan a sequence of actions. However, such models are often learned from noisy data, and deploying them in the real world might lead an agent to catastrophic states. Therefore, methods in the Safe-RL literature focus on incorporating safety-related constraints into the training process to mitigate such issues.
While Safety in RL is a vast field in and of itself (?), we consider two specific categories in this work: Safe Learning with constraints and Safe Exploration. The former subjects the learning process to one or more constraints $c_{i}\in C$ (?). Depending on the necessity of strictness, these can be incorporated in many different ways, such as safety in expectation, safety in values, safe trajectories, and safe states and actions. We can formulate this as
$\max_{\pi\in\Pi}\mathbb{E}_{\pi}(G)\,\,\,s.t.\,\,\,c_{i}=\{h_{i}\leq\alpha\},$ (6)
where $h_{i}$ can be a function related to the returns, trajectories, values, states, and actions, and $\alpha$ is a safety threshold. Consequently, side information can be used in the formulation of such constraints.
On the other hand, Safe Exploration modifies the exploration process subject to external knowledge, which in our case translates to incorporating side information into the exploration process. While intuitively, this overlaps with the usage of side information for directed exploration, a distinguishing feature of this work is the final goal of this directed exploration, which is to be safe, which might come at the cost of sample efficiency and/or generalization.
@@ 5 Structure as Side Information
Structure can be considered a particular kind of side information about decomposability. In this section, we discuss the nature of decomposability and the various ways it can be represented. In Section 5.1, we explain the impact of structural side information by explaining how it decomposes complex systems and categorizes such decompositions into four archetypes. In Section 5.2 - Section 5.5, we discuss these archetypes further to connect them with existing literature.
@@ 5.1 Decomposability and Structural Archetypes
Decomposability is the property of a system that allows breaking it down into smaller components or subsystems that can be analyzed, understood, and potentially learned more efficiently than the larger system, independently (?). In a decomposable system, the short-term behavior of each subsystem is approximately independent of the short-term behavior of the other subsystems. In the long run, the behavior of any subsystem depends on the behavior of the other subsystems only in an aggregated way.
Concerning the RL pipeline in Figure 2, we can see decomposability along two axes: (i) Problem Decompositioni.e., the environment parameterization, states, actions, transitions, and rewards; (ii) Solution Decompositioni.e., the learned policies, value functions, models, and training procedures. The spectrum of decomposability (?) provides an intuitive way to understand where a system lies in this regard. On one end of the spectrum, problems are non-decomposable, while on the other end, problems can be decomposed into weakly interacting sub-problems. Similarly, solutions on the former are monolithic, while those on the latter are modular. We capture this problem-solution interplay by marking four different archetypes of decomposability. Decomposition can be incorporated by learning appropriate representations at the granularity of the decomposition, as shown in Figure 3. Thus, the following sections use the terms decompositions and representations interchangeably since the representations that we are referring to in this work particularly target decompositions.
Figure 3: Spectrum of Decomposability and Structural Archetypes. On the left end of the spectrum exist monolithic structural decompositions where a latent representation of $\mathcal{X}$ can be learned and incorporated as an inductive bias. Moving towards the right, we can learn multiple latent representations, albeit in a monolithic solution. These are factored representations. Further ahead, we see the emergence of interactionally complex decompositions, where knowledge about factorization and how they relate to each other might be essential and can be incorporated using relational representations. Finally, we see fully distributed subsystems that can be learned using individual policies. We call these modular representations.
@@ 5.2 Latent Decomposition
Latent representations can be helpful in complex environments where the underlying structure is unclear or non-stationary. Under this view, a pipeline component $\mathcal{X}$ can be approximated by a latent representation $\kappa$, which can then be integrated into the learning process. The quantities in the learning process that depend on $\mathcal{X}$ can now be re-conditioned on $\kappa$:
$Z:\mathcal{X}\to\kappa.$ (7)
@@
Latent representations of states are used for tackling scenarios such as rich observation spaces (?) and contextual settings (?). Latent actions have been similarly explored in settings with stochastic action sets (?).
@@
While latent states allow decomposing transition matrices, another way to approach the problem directly is to decompose transition matrices into low-rank approximations. Linear MDPs (?) and applications in Model-based RL (?) have studied this form of direct decomposition. A similar decomposition can also be applied to rewards by assuming the reward signal to be generated from a latent function that can be learned as an auxiliary learning objective (?).
@@ 5.3 Factored Decomposition
The factored decomposition moves slightly away from the monolithic nature by representing $\mathcal{X}$ using (latent) factors $\mathcal{K}=\{\kappa_{1},\dots,\kappa_{N}$. A crucial aspect of factorization is that the factors can potentially impose some form of conditional independence in their effects on the learning dynamics.
$Z:\mathcal{X}\to\mathcal{K}$ (8)
@@
Factored state and action spaces have been explored in the Factored MDPs (?, ?, ?). Methods in this setting traditionally capture subsequent state distribution using mechanisms such as Dynamic Bayesian Networks (?).
Factored action representations have also been used for tackling high-dimensional actions (?). These methods either impose a factorized structure on subsets of a high-dimensional action set (?) or impose this structure through the Q-values that lead to the final action (?). Crucially, these methods can potentially exploit some form of independence resulting from such factorization, either in the state representations or transitions.
@@
Combined with factored states or modeled independently, factored rewards have been used to model perturbed rewards (?), or to factor latent variable models using causal priors (?). While Factored MDPs do not naturally lead to factored policies, combining state and reward factorization can lead to factorization of value functions (?, ?).
@@ 5.4 Relational Decomposition
In addition to representing the problem using a set of factors, we can also utilize information about how different factors interact. Usually, these relations exist between entities in a scene and are used to formulate learning methods based on inductive logic (?). Traditionally, such relations were limited to first-order logic, but the relational structure has also been captured through graphs. Mathematically, side information $Z$ takes the original entity $\mathcal{X}$ as an input and maps it to to a function $\phi$ over the set of factors $\mathcal{K}=\{\kappa_{i},\dots\kappa_{N}\}$. Therefore, $Z$ now outputs functions over $\mathcal{K}$.
$Z:\mathcal{X}\to\phi\big{(}\mathcal{K}\big{)}$ (9)
Specifically, $\phi$ is a function that describes relations between groups of $m$ entities ($m$ being the order of the relation) in $\mathcal{K}$. Thus, the inputs to $\phi$ are $m$-tuples of factors, which it maps to a multiset of symbols $\{\psi\}^{+}$ (such as coordinates, distance measures, or logical predicates). In other words, $\phi$ describes relations between $m$ input entities using symbols $\psi$, and the multiset allows us to more generally describe the case where similar relations can exist between different entities:
$\phi:(\mathcal{K})^{m}\to\{\psi\}^{+}.$ (10)
Such representations allow talking about generalization over the relationship between entities in $\mathcal{K}$ and different forms of $\phi$ and help us circumvent the dimensionality of enumerative spaces.
Relational States and Actions.
Classically, relational representations have been used to model state spaces in Relational MDPs (?, ?) and Object-Oriented MDPs (?). They represent factored state spaces using first-order representations of objects, predicates, and functions to describe a set of ground MDPs. Such representations can capture interactionally complex structures between entities much more efficiently. Additionally, permutations of the interactions between the entities can help define new MDPs that differ in their dynamics, thus contributing towards work in generalization.
@@
Action relations help tackle instances where the agent has multiple possible actions available and the set of actions is significantly large. These methods capture relations using either attention mechanisms (?, ?) or graphs (?), thus offering scalability to high-dimensional action spaces. Additionally, relations between states and actions have helped define notions such as intents and affordances (?, ?).
Relational Value Functions and Policies.
Traditional work in Relational MDPs has also explored ways to represent and construct first-order representations of value functions and/or policies to generalize to new instances. These include Regression Trees (?), Decision Lists (?), Algebraic Decision Diagrams (?), Linear Basis Functions (?, ?), and Graph Laplacians (?). Recent approaches have started looking into DNN representations (?, ?), with extensions into modeling problem aspects such as morphology in Robotic tasks (?) in a relational manner, or learning diffusion operators for constructing intrinsic rewards (?).
Relational Tasks.